OpenStudy (anonymous):

Solve S = 4v2 for v ?

mathslover (mathslover):

did u mean this ? \[\huge{s=4v^2}\]

mathslover (mathslover):

if u meant this then v can be calculated by rooting both sides i meant this : \[\large{\sqrt{s}=\sqrt{4v^2}}\]

mathslover (mathslover):

can u solve this now @alexawb

OpenStudy (anonymous):

thank you, although not really sure how to find \[\sqrt{4v ^{2}}\] ?

mathslover (mathslover):

or should i give u more hint ?

OpenStudy (anonymous):

yes please :/

mathslover (mathslover):

\[\sqrt{ab}=\sqrt{a}*\sqrt{b}\]

mathslover (mathslover):

\[\huge{\sqrt{a}*\sqrt{b}}\] can u do this now ?

mathslover (mathslover):

therefore \[\huge{\sqrt{4v^2}=\sqrt{4}*\sqrt{v^2}}\] \[\huge{\sqrt{4v^2}=2v}\]

OpenStudy (anonymous):

ohhhh. awesome, thank you so much

OpenStudy (anonymous):

i see now :)

mathslover (mathslover):

my pleasure ... great to hear that u got this question